If U = I + 3j - 2k and V = 4i - 2j - 4k are vectors, find 3U + V

10
14
4
11
1
Expert's answer
2014-09-23T09:54:40-0400
3U + V=3(i + 3j - 2k)+(4i - 2j - 4k)=3i+ 9j - 6k+4i - 2j - 4k=7i+7j-10k